I am trying to help my son with this problem and I can't figure out where to begin on this one. I was fine until fractions came into play
2/3 +13y+1/4=2
Should he add the fractions first and then take next steps?

Answer by scott8148(6628)

not being a big fan of fractions, myself ; I clear them by multiplying by the LCD (in this case , 12)

8 + 156y + 3 = 24

156y = 13

y = 13/156 = 1/12